# Variational Diffusion Posterior Sampling with Midpoint Guidance

### v1

@@ 4 Experiments
An electrocardiogram (ECG) is an electrical recording of the heart’s activity in which the signals generated by the heartbeats are recorded in order to diagnose various cardiac conditions such as cardiac arrhythmias and conduction abnormalities. Unlike static images, ECGs constitute complex time-series data consisting of 12 electrical signals acquired using 10 electrodes, 4 of which are attached to the limbs and record the ’limb leads’, while the others record the ’precordial leads’ around the heart. We study two conditional generation problems in ECGs. The first is a forecasting or missing-block (MB) reconstruction problem, where one half of the ECG is reconstructed from the other (see fig. 7). This task evaluates the algorithm’s ability to capture temporal information to predict a coherent signal. The second problem is an inpainting or missing-leads (ML) reconstruction, where the entire ECG is reconstructed from the lead I (see fig. 5). The question is whether we can capture the subtle information contained in lead I and reconstruct a coherent ECG with the same diagnosis as the real ECG. This task is challenging because lead I, being acquired with limb electrodes far from the heart, may contain very subtle features related to specific cardiac conditions. We train a state-space diffusion model Goel et al. (2022) to generate ECGs using the 20k training ECGs from the ptb-xl dataset Wagner et al. (2020), and benchmark the posterior sampling algorithm on the 2k test ECGs (see section D.5). We report the Mean Absolute Error (MAE) and the Root Mean Squared Error (RMSE) between ground-truth and reconstructions in Table 4. We demonstrate that a diffusion model trained to generate ECGs can serve as a prior to solve imputation tasks without additional training or fine-tuning, yielding superior results to a diffusion model trained conditionally on the MB task Alcaraz & Strodthoff (2022). The rationale for this result is discussed in Section D.5.2. We report in Section 4 the balanced accuracy of diagnosing Left Bundle Branch Block (LBBB), Right Bundle Branch Block (RBBB), Atrial Fibrillation (AF), and Sinus Bradycardia (SB) using the downstream classifier proposed in Strodthoff et al. (2020) (see section D.5.1), applied to both ground-truth and to reconstructed samples from lead I. See section D.5.2 for sample figures. MGPS outperforms all other posterior sampling algorithms with just 50 diffusion steps.
@@ 5 Conclusion
Although the proposed method is promising, it has certain limitations that invite further exploration. First, a detailed analysis of how the algorithm’s performance depends on the choice of the intermediate time steps $({\ell_{k}})_{k=1}^{n-1}$ remains a challenging but critical subject for future research. Although our image and ECG experiments suggest that setting ${\ell_{k}}=\lfloor k/2\rfloor$ leads to significant performance gains on most tasks, we have observed that using an adaptive sequence ${\ell_{k}}=\lfloor\eta_{k}k\rfloor$, where $\eta_{k}$ increases as $k$ decreases, further enhances results on certain tasks, such as Gaussian and motion deblurring, as well as when using latent diffusion models. A second direction for improvement lies in refining the approximation of $g_{k}$ which we believe could lead to overall algorithmic improvements. Specifically, devising an approximation $\hat{g}^{\theta}_{k}$ such that $\nabla\log\hat{g}^{\theta}_{k}$ does not require a vector-Jacobian product could significantly reduce the algorithm’s runtime. Lastly, we see potential in using the two-stage approximation introduced in our warm-start strategy (see Algorithm 3) at every diffusion step. Although this technique is promising, it currently leads to instability as $k$ decreases. Understanding and resolving this instability will be a key focus of our future work.
@@ A.1 Denoising diffusion probabilistic models
DDPMs define generative models for $q$ relying only on parametric approximations $(\smash{m^{\theta}_{0|t}})_{t=1}^{T}$ of the denoisers $(\smash{m_{0|t}})_{t=1}^{T}$. These approximate denoisers are usually defined through the parameterization
$\smash{m^{\theta}_{0|t}}(x_{t})=(x_{t}-\sqrt{1-\alpha_{t}}\smash{\epsilon^{\theta}_{t}}(x_{t}))/\sqrt{\alpha_{t}}$ (A.1)
@@
$\sum_{t=1}^{T}w_{t}\mathbb{E}\left[\|\epsilon_{t}-\smash{\epsilon^{\theta}_{t}}(\sqrt{\alpha_{t}}X_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t})\|^{2}\right]\,,$ (A.2)
w.r.t. the neural network parameter $\theta$, where $(\epsilon_{t})_{t=1}^{T}$ are i.i.d. standard normal vectors, $X_{0}\sim q$, and $(w_{t})_{t=1}^{T}$ are some nonnegative weights. Having trained the denoisers, the generative model for $q$ is defined as follows. Let $(t_{k})_{k=0}^{n}$ be an increasing sequence of time steps in $\llbracket 0,T\rrbracket$ with $t_{0}=0$. We assume that $t_{n}$ is large enough so that $q_{t_{n}}(x_{t_{n}})=\int q(x_{0})q_{t_{n}|0}(x_{t_{n}}|x_{0})\,\mathrm{d}x_{0}$ is approximately the density of a multivariate standard normal distribution. For convenience we assign the index $k$ to any quantity depending on $t_{k}$; e.g., we denote $q_{t_{k}}$ by $q_{k}$. Consider the backward decomposition
$q_{0:n}(x_{0:n})=q(x_{0})\prod_{k=0}^{n-1}q_{k+1|k}(x_{k+1}|x_{k})=q_{n}(x_{n})\prod_{k=0}^{n-1}q_{k|k+1}(x_{k}|x_{k+1})$
of the forward process initialized at $q$, where $q_{k|k+1}(x_{k}|x_{k+1})\propto q_{k}(x_{k})q_{k+1|k}(x_{k+1}|x_{k})$. Next, for $(j,\ell,k)\in\llbracket 0,n\rrbracket^{3}$ such that $j<\ell<k$, define
$m_{\ell|j,k}(x_{j},x_{k})$ :$\vcentcolon=\frac{\sqrt{\alpha_{\ell}{}/\alpha_{j}}(1-\alpha_{k}/\alpha_{\ell})}{1-\alpha_{k}/\alpha_{j}}x_{j}+\frac{\sqrt{\alpha_{k}/\alpha_{\ell}}(1-\alpha_{\ell}/\alpha_{j})}{1-\alpha_{k}/\alpha_{j}}x_{k}\,,$ (A.3)
@@
Then the bridge kernel
$q_{\ell|j,k}(x_{\ell}|x_{j},x_{k})=q_{\ell|j}(x_{\ell}|x_{j})q_{k|\ell}(x_{k}|x_{\ell})\big{/}q_{k|j}(x_{k}|x_{j})$ (A.5)
is the density of a Gaussian distribution with mean $m_{\ell|j,k}(x_{j},x_{k})$ and covariance $v_{\ell|j,k}\mathrm{I}_{d}$. Using the bridge kernel, $q_{0:n}$ is approximated using the variational approximation
$p^{\theta}_{0:n}(x_{0:n})=p^{\theta}_{n}(x_{n})\prod_{k=0}^{n-1}p^{\theta}_{k|k+1}(x_{k}|x_{k+1})\,,$
where $p^{\theta}_{0|1}(x_{0}|x_{1})=\mathcal{N}(x_{0};\smash{m^{\theta}_{0|1}}(x_{1}),v_{0|1}\mathrm{I}_{d})$, $v_{0|1}$ being a tunable parameter, and
$p^{\theta}_{k|k+1}(x_{k}|x_{k+1})\vcentcolon=q_{k|0,k+1}(x_{k}|\smash{m^{\theta}_{0|k+1}}(x_{k+1}),x_{k+1})\,\quad k\in\llbracket 1,n-1\rrbracket\,.$ (A.6)
@@ A.2 Latent diffusion models
Latent diffusion models (LDM) Rombach et al. (2022) define a DDM in a latent space. Let $\mathcal{E}:\mathbb{R}^{d}\to\mathbb{R}^{p}$ be an encoder function and $\mathcal{D}:\mathbb{R}^{p}\to\mathbb{R}^{d}$ a decoder function. We assume that these functions satisfy $\mathcal{D}_{\sharp}\mathcal{E}_{\sharp}q\approx q$, where, for instance, $\mathcal{E}_{\sharp}q$ denotes the law of the random variable $\mathcal{E}(X)$ where $X\sim q$. A LDM approximating $q$ is given by $\mathcal{D}_{\sharp}p^{\theta}_{0}$, where $p^{\theta}_{0}$ is a diffusion model trained on samples from $\mathcal{E}_{\sharp}q$. Finally, when solving inverse problems with LDMs, we assume that the target distribution is instead
$\pi(x)\propto g(x)\mathcal{D}_{\sharp}\mathcal{E}_{\sharp}q(x)\,.$
Let $X\sim\pi$, $D\sim\mathcal{D}_{\sharp}\mathcal{E}_{\sharp}q$, $E\sim\mathcal{E}_{\sharp}q$, and $Z\sim\overline{\pi}$, where $\overline{\pi}(z)\propto g(\mathcal{D}_{\sharp}(z))\mathcal{E}_{\sharp}q(z)$. For any bounded function $f$ on $\mathbb{R}^{d}$, we have, following the definition of $\pi$, that
$\mathbb{E}[f(X)]=\frac{\mathbb{E}[g(D)f(D)]}{\mathbb{E}[g(D)]}=\frac{\mathbb{E}[g(\mathcal{D}(E))f(\mathcal{D}(E))]}{\mathbb{E}[g(\mathcal{D}(E))]}=\mathbb{E}[f(\mathcal{D}(Z))]\,.$
Hence $\mathsf{Law}(X)=\mathsf{Law}(\mathcal{D}(Z))$. As a result, to sample approximately from $\pi$, we first sample approximately from $\overline{\pi}$ using any diffusion posterior sampling algorithm with pre-trained DDM for $\mathcal{E}_{\sharp}q$, then decode the obtained samples using $\mathcal{D}$.
@@ A.3 Midpoint decomposition
$\smash{\pi}_{i}(x_{i})q_{j|i}(x_{j}|x_{i})=\smash{\pi}_{j}(x_{j})\smash{\pi}_{i|j}(x_{i}|x_{j})\,,$ (A.7)
where $\smash{\pi}_{i|j}(x_{i}|x_{j})\vcentcolon=\pi_{i}(x_{i})q_{j|i}(x_{j}|x_{i})/\pi_{j}(x_{j})$ and integrates to one.
@@
$q_{k|\ell,k+1}(x_{k}|x_{\ell},x_{k+1})$ $=\frac{q_{k|\ell}(x_{k}|x_{\ell})q_{k+1|k}(x_{k+1}|x_{k})}{q_{k+1|\ell}(x_{k+1}|x_{\ell})}$
$=\frac{\smash{\pi}_{\ell}(x_{\ell})q_{k|\ell}(x_{k}|x_{\ell})q_{k+1|k}(x_{k+1}|x_{k})}{\smash{\pi}_{\ell}(x_{\ell})q_{k+1|\ell}(x_{k+1}|x_{\ell})}$
$=\frac{\smash{\pi}_{\ell|k}(x_{\ell}|x_{k})\smash{\pi}_{k|k+1}(x_{k}|x_{k+1})\smash{\pi}_{k+1}(x_{k+1})}{\smash{\pi}_{\ell|k+1}(x_{\ell}|x_{k+1})\smash{\pi}_{k+1}(x_{k+1})}$
$=\frac{\smash{\pi}_{\ell|k}(x_{\ell}|x_{k})\smash{\pi}_{k|k+1}(x_{k}|x_{k+1})}{\smash{\pi}_{\ell|k+1}(x_{\ell}|x_{k+1})}\enspace.$
@@
$\smash{\pi}_{k|k+1}(x_{k}|x_{k+1})$ $=\int\smash{\pi}_{\ell|k}(x_{\ell}|x_{k})\smash{\pi}_{k|k+1}(x_{k}|x_{k+1})\,\mathrm{d}x_{\ell}$
$=\int q_{k|\ell,k+1}(x_{k}|x_{\ell},x_{k+1})\smash{\pi}_{\ell|k+1}(x_{\ell}|x_{k+1})\,\mathrm{d}x_{\ell}\,.$

### v2

@@ 4 Experiments
We now explore posterior sampling algorithms beyond image tasks to show their versatility and public health impact. Cardiovascular diseases cause one-third of global deaths, and better detection can improve management. Wearables like smartwatches can enhance diagnosis by capturing brief symptom episodes, particularly for conditions like paroxysmal atrial fibrillation (AF), which may go undetected during routine medical visits. However, they offer only a partial ECG view (Lead I instead of 12 leads) and a recent study showed the Apple Watch detected AF in only 34 of 90 episodes Seshadri et al. (2020). To address this limitation, we propose using posterior sampling algorithms to reconstruct incomplete electrocardiograms (ECG). An ECG is an electrical recording of the heart’s activity in which the signals generated by the heartbeats are recorded in order to diagnose various cardiac conditions such as cardiac arrhythmias and conduction abnormalities. Unlike static images, ECGs constitute complex time-series data consisting of 12 electrical signals acquired using 10 electrodes, 4 of which are attached to the limbs and record the ’limb leads’, while the others record the ’precordial leads’ around the heart. We study two conditional generation problems in ECGs. The first is a forecasting or missing-block (MB) reconstruction problem, where one half of the ECG is reconstructed from the other (see fig. 8). This task evaluates the algorithm’s ability to capture temporal information to predict a coherent signal. The second problem is an inpainting or missing-leads (ML) reconstruction, where the entire ECG is reconstructed from the lead I (see fig. 6). The question is whether we can capture the subtle information contained in lead I and reconstruct a coherent ECG with the same diagnosis as the real ECG. This task is challenging because lead I, being acquired with limb electrodes far from the heart, may contain very subtle features related to specific cardiac conditions. We train a state-space diffusion model Goel et al. (2022) to generate ECGs using the 20k training ECGs from the ptb-xl dataset Wagner et al. (2020), and benchmark the posterior sampling algorithm on the 2k test ECGs (see section D.5). We report the Mean Absolute Error (MAE) and the Root Mean Squared Error (RMSE) between ground-truth and reconstructions in Table 4. We demonstrate that a diffusion model trained to generate ECGs can serve as a prior to solve imputation tasks without additional training or fine-tuning, yielding superior results to a diffusion model trained conditionally on the MB task Alcaraz & Strodthoff (2022). The rationale for this result is discussed in Section D.5.2. We report in Section 4 the balanced accuracy of diagnosing Left Bundle Branch Block (LBBB), Right Bundle Branch Block (RBBB), Atrial Fibrillation (AF), and Sinus Bradycardia (SB) using the downstream classifier proposed in Strodthoff et al. (2020) (see section D.5.1), applied to both ground-truth and to reconstructed samples from lead I. See section D.5.2 for sample figures. MGPS outperforms all other posterior sampling algorithms with just 50 diffusion steps.
@@ 5 Conclusion
Although the proposed method is promising, it has certain limitations that invite further exploration. First, a detailed analysis of how the algorithm’s performance depends on the choice of the intermediate time steps $({\ell_{k}})_{k=1}^{n-1}$ remains a challenging but critical subject for future research. Although our image and ECG experiments suggest that setting ${\ell_{k}}=\lfloor k/2\rfloor$ leads to significant performance gains on most tasks, we have observed that using an adaptive sequence ${\ell_{k}}=\lfloor\eta_{k}k\rfloor$, where $\eta_{k}$ increases as $k$ decreases, further enhances results on certain tasks, such as Gaussian and motion deblurring, as well as when using latent diffusion models. A second direction for improvement lies in refining the approximation of $p_{k}({\boldsymbol{y}}|\cdot)$ which we believe could lead to overall algorithmic improvements. Specifically, devising an approximation $\hat{p}^{\theta}_{k}({\boldsymbol{y}}|\cdot)$ such that $\nabla\log\hat{p}^{\theta}_{k}({\boldsymbol{y}}|\cdot)$ does not require a vector-Jacobian product could significantly reduce the algorithm’s runtime. Lastly, we see potential in using the two-stage approximation introduced in our warm-start strategy (see Algorithm 3) at every diffusion step. Although this technique is promising, it currently leads to instability as $k$ decreases. Understanding and resolving this instability will be a key focus of our future work.
@@ A.1 Denoising diffusion probabilistic models
DDPMs define generative models for $q$ relying only on parametric approximations $(\smash{\boldsymbol{m}^{\theta}_{0|t}})_{t=1}^{T}$ of the denoisers $(\smash{\boldsymbol{m}_{0|t}})_{t=1}^{T}$. These approximate denoisers are usually defined through the parameterization
$\smash{\boldsymbol{m}^{\theta}_{0|t}}(\boldsymbol{x}_{t})=(\boldsymbol{x}_{t}-\sqrt{1-\alpha_{t}}\smash{\boldsymbol{\epsilon}^{\theta}_{t}}(\boldsymbol{x}_{t}))/\sqrt{\alpha_{t}}$ (A.1)
@@
$\sum_{t=1}^{T}w_{t}\mathbb{E}\left[\|\boldsymbol{\epsilon}_{t}-\smash{\boldsymbol{\epsilon}^{\theta}_{t}}(\sqrt{\alpha_{t}}\mathbf{X}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{t})\|^{2}\right]\,,$ (A.2)
w.r.t. the neural network parameter $\theta$, where $(\boldsymbol{\epsilon}_{t})_{t=1}^{T}$ are i.i.d. standard normal vectors, $\mathbf{X}_{0}\sim q$, and $(w_{t})_{t=1}^{T}$ are some nonnegative weights. Having trained the denoisers, the generative model for $q$ is defined as follows. Let $(t_{k})_{k=0}^{n}$ be an increasing sequence of time steps in $\llbracket 0,T\rrbracket$ with $t_{0}=0$. We assume that $t_{n}$ is large enough so that $q_{t_{n}}(\boldsymbol{x}_{t_{n}})=\int q(\boldsymbol{x}_{0})q_{t_{n}|0}(\boldsymbol{x}_{t_{n}}|\boldsymbol{x}_{0})\,\mathrm{d}\boldsymbol{x}_{0}$ is approximately the density of a multivariate standard normal distribution. For convenience we assign the index $k$ to any quantity depending on $t_{k}$; e.g., we denote $q_{t_{k}}$ by $q_{k}$. Consider the backward decomposition
$q_{0:n}(\boldsymbol{x}_{0:n})=q(\boldsymbol{x}_{0})\prod_{k=0}^{n-1}q_{k+1|k}(\boldsymbol{x}_{k+1}|\boldsymbol{x}_{k})=q_{n}(\boldsymbol{x}_{n})\prod_{k=0}^{n-1}q_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})$
of the forward process initialized at $q$, where $q_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})\propto q_{k}(\boldsymbol{x}_{k})q_{k+1|k}(\boldsymbol{x}_{k+1}|\boldsymbol{x}_{k})$. Next, for $(j,\ell,k)\in\llbracket 0,n\rrbracket^{3}$ such that $j<\ell<k$, define
$\boldsymbol{m}_{\ell|j,k}(\boldsymbol{x}_{j},\boldsymbol{x}_{k})$ :$\vcentcolon=\frac{\sqrt{\alpha_{\ell}{}/\alpha_{j}}(1-\alpha_{k}/\alpha_{\ell})}{1-\alpha_{k}/\alpha_{j}}\boldsymbol{x}_{j}+\frac{\sqrt{\alpha_{k}/\alpha_{\ell}}(1-\alpha_{\ell}/\alpha_{j})}{1-\alpha_{k}/\alpha_{j}}\boldsymbol{x}_{k}\,,$ (A.3)
@@
Then the bridge kernel is
$q_{\ell|j,k}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{j},\boldsymbol{x}_{k})$ $=q_{\ell|j}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{j})q_{k|\ell}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell})\big{/}q_{k|j}(\boldsymbol{x}_{k}|\boldsymbol{x}_{j})$
$=\mathcal{N}(\boldsymbol{x}_{\ell};\boldsymbol{m}_{\ell|j,k}(\boldsymbol{x}_{j},\boldsymbol{x}_{k}),v_{\ell|j,k}\mathbf{I}_{d})\,.$ (A.5)
Using the bridge kernel, $q_{0:n}$ is approximated using the variational approximation
$p^{\theta}_{0:n}(\boldsymbol{x}_{0:n})=p^{\theta}_{n}(\boldsymbol{x}_{n})\prod_{k=0}^{n-1}p^{\theta}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})\,,$
where $p^{\theta}_{0|1}(\boldsymbol{x}_{0}|\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}_{0};\smash{\boldsymbol{m}^{\theta}_{0|1}}(\boldsymbol{x}_{1}),v_{0|1}\mathbf{I}_{d})$, $v_{0|1}$ being a tunable parameter, and
$p^{\theta}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})\vcentcolon=q_{k|0,k+1}(\boldsymbol{x}_{k}|\smash{\boldsymbol{m}^{\theta}_{0|k+1}}(\boldsymbol{x}_{k+1}),\boldsymbol{x}_{k+1})\,\quad k\in\llbracket 1,n-1\rrbracket\,.$ (A.6)
@@ A.2 Latent diffusion models
Latent diffusion models (LDM) Rombach et al. (2022) define a DDM in a latent space. Let $\mathcal{E}:\mathbb{R}^{d}\to\mathbb{R}^{p}$ be an encoder function and $\mathcal{D}:\mathbb{R}^{p}\to\mathbb{R}^{d}$ a decoder function. We assume that these functions satisfy $\mathcal{D}_{\sharp}\mathcal{E}_{\sharp}q\approx q$, where, for instance, $\mathcal{E}_{\sharp}q$ denotes the law of the random variable $\mathcal{E}(\mathbf{X})$ where $\mathbf{X}\sim q$. A LDM approximating $q$ is given by $\mathcal{D}_{\sharp}p^{\theta}_{0}$, where $p^{\theta}_{0}$ is a diffusion model trained on samples from $\mathcal{E}_{\sharp}q$. Finally, when solving inverse problems with LDMs, we assume that the target distribution is instead
$\pi(\boldsymbol{x})\propto p({\boldsymbol{y}}|\boldsymbol{x})\mathcal{D}_{\sharp}\mathcal{E}_{\sharp}q(\boldsymbol{x})\,.$
Let $\mathbf{X}\sim\pi$, $D\sim\mathcal{D}_{\sharp}\mathcal{E}_{\sharp}q$, $E\sim\mathcal{E}_{\sharp}q$, and $\mathbf{Z}\sim\overline{\pi}$, where $\overline{\pi}(\boldsymbol{z})\propto p({\boldsymbol{y}}|\mathcal{D}_{\sharp}(\boldsymbol{z}))\mathcal{E}_{\sharp}q(\boldsymbol{z})$. For any bounded function $f$ on $\mathbb{R}^{d}$, we have, following the definition of $\pi$, that
$\mathbb{E}[f(\mathbf{X})]=\frac{\mathbb{E}[p({\boldsymbol{y}}|D)f(D)]}{\mathbb{E}[p({\boldsymbol{y}}|D)]}=\frac{\mathbb{E}[p({\boldsymbol{y}}|\mathcal{D}(E))f(\mathcal{D}(E))]}{\mathbb{E}[p({\boldsymbol{y}}|\mathcal{D}(E))]}=\mathbb{E}[f(\mathcal{D}(\mathbf{Z}))]\,.$
Hence $\mathsf{Law}(\mathbf{X})=\mathsf{Law}(\mathcal{D}(\mathbf{Z}))$. As a result, to sample approximately from $\pi$, we first sample approximately from $\overline{\pi}$ using any diffusion posterior sampling algorithm with pre-trained DDM for $\mathcal{E}_{\sharp}q$, then decode the obtained samples using $\mathcal{D}$.
@@ A.3 Midpoint decomposition
$\smash{\pi}_{i}(\boldsymbol{x}_{i})q_{j|i}(\boldsymbol{x}_{j}|\boldsymbol{x}_{i})=\smash{\pi}_{j}(\boldsymbol{x}_{j})\smash{\pi}_{i|j}(\boldsymbol{x}_{i}|\boldsymbol{x}_{j})\,,$ (A.7)
where $\smash{\pi}_{i|j}(\boldsymbol{x}_{i}|\boldsymbol{x}_{j})\vcentcolon=\pi_{i}(\boldsymbol{x}_{i})q_{j|i}(\boldsymbol{x}_{j}|\boldsymbol{x}_{i})/\pi_{j}(\boldsymbol{x}_{j})$ and integrates to one.
@@
$q_{k|\ell,k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell},\boldsymbol{x}_{k+1})$ $=\frac{q_{k|\ell}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell})q_{k+1|k}(\boldsymbol{x}_{k+1}|\boldsymbol{x}_{k})}{q_{k+1|\ell}(\boldsymbol{x}_{k+1}|\boldsymbol{x}_{\ell})}$
$=\frac{\smash{\pi}_{\ell}(\boldsymbol{x}_{\ell})q_{k|\ell}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell})q_{k+1|k}(\boldsymbol{x}_{k+1}|\boldsymbol{x}_{k})}{\smash{\pi}_{\ell}(\boldsymbol{x}_{\ell})q_{k+1|\ell}(\boldsymbol{x}_{k+1}|\boldsymbol{x}_{\ell})}$
$=\frac{\smash{\pi}_{\ell|k}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{k})\smash{\pi}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})\smash{\pi}_{k+1}(\boldsymbol{x}_{k+1})}{\smash{\pi}_{\ell|k+1}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{k+1})\smash{\pi}_{k+1}(\boldsymbol{x}_{k+1})}$
$=\frac{\smash{\pi}_{\ell|k}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{k})\smash{\pi}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})}{\smash{\pi}_{\ell|k+1}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{k+1})}\enspace.$
@@
$\smash{\pi}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})$ $=\int\smash{\pi}_{\ell|k}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{k})\smash{\pi}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})\,\mathrm{d}\boldsymbol{x}_{\ell}$
$=\int q_{k|\ell,k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{\ell},\boldsymbol{x}_{k+1})\smash{\pi}_{\ell|k+1}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{k+1})\,\mathrm{d}\boldsymbol{x}_{\ell}\,.$